OpenStudy (konradzuse):

Integration by parts. integral of xe^(-5x) dx

OpenStudy (konradzuse):

u = x du = dx dv = e^(-5x) v = (-1/5)x e^(-5x)?

OpenStudy (anonymous):

I'm not familiar with your notation, but I think you're on the right track.

OpenStudy (anonymous):

The standard way is use the formula for integration by parts. integral u(x)v(x) dx = u(x) integral v(x)dx - integral [ ( integral v(x)dx ) (du/dx) ] dx Let u(x) = x and v(x) = e^5x. However you can often do these quicker by guessing what part of the answer will be. Knowing that the derivative of e^x is e^x, and knowing how to differentiate the product of two functions, you can guess the answer will probably contain the term like xe^5x. d/dx (xe^5x) = 5xe^5x + e^5x So integral xe^5x dx = 1/5 integral (5xe^5x + e^5x - e^5x) dx = 1/5 integral (5xe^5x + e^5x) dx - 1/5 integral e^5x dx = 1/5 xe^5x - 1/25 e^5x + C

OpenStudy (anonymous):

@Thomas9 & others, for reference, Integration by Parts: \[\int\limits_{a}^{b}U \cdot dV = U \cdot V - \int\limits_{a}^{b} V \cdot dU\]

OpenStudy (konradzuse):

^yup

OpenStudy (konradzuse):

so getting back to my solving its u * v - int vdu \[x * e^{-5x} - -\frac{1}{5}\int\limits e^{-5x} dx\]

OpenStudy (konradzuse):

\[x * e^{-5x} + \frac{1}{25} e^{-5x} + c\]

OpenStudy (anonymous):

You're missing -1/5 in the first term.

OpenStudy (konradzuse):

oh woops yeah.

OpenStudy (konradzuse):

\[-\frac{1}{5}x * e^{-5x} + \frac{1}{25}e^{-5x} + c?\]

OpenStudy (anonymous):

The second + is a minus.

OpenStudy (anonymous):

first plus actually.

OpenStudy (konradzuse):

wasn't it - -1/5 intgral?

OpenStudy (konradzuse):

since initially it was a - the integral, and the dx = -1/5du?

OpenStudy (anonymous):

You get another -1/5 from integrating in the last step.

OpenStudy (konradzuse):

oh yeah tha't sright....

OpenStudy (konradzuse):

forgot about that :). ty